OpenStudy (anonymous):

I need help finding an equation of the line with the slope=-2, containing the points (3,-1).

OpenStudy (anonymous):

y=mx+b, plug in the values for x, y and the slope to find b

OpenStudy (phi):

one form is y - y1= m( x- x1) where m is the slope and (x1,y1) is a point on the line

OpenStudy (anonymous):

-1=(-2)(3)+b -1=-6+b b=5 y=-2x+5

OpenStudy (phi):

Here is a list of different forms of an equation for a line http://en.wikipedia.org/wiki/Linear_equation#Forms_for_2D_linear_equations I gave you point-slope emilie showed you slope-intercept

OpenStudy (anonymous):

thank you.